# GraphOracle: A Foundation Model for Knowledge Graph Reasoning

### v1

@@ 6 Conclusion
In this work, we introduced GraphOracle, a relation-centric foundation model for unifying reasoning across heterogeneous KGs. By converting KGs into relation-dependency graphs, our approach explicitly encodes compositional patterns among relations, yielding domain-invariant embeddings. Experiments on 31 diverse benchmarks showed consistent state-of-the-art performance, improving mean reciprocal rank by up to 35% over baselines with minimal adaptation. Pre-training on diverse KGs followed by brief fine-tuning enables effective zero-shot and few-shot generalization to unseen entities, relations, and entire graphs. We also demonstrated that GraphOracle’s performance can be enhanced by integrating external information through GraphOracle+, which leverages foundation model embeddings for improved initialization. Ablation studies confirmed the essential contributions of the relation-dependency graph and multi-head attention components. These findings establish relation-dependency pre-training as a scalable approach toward universal KG reasoning, opening new avenues for cross-domain applications.
Appendix A Limitations and Future Work
@@
Appendix B Comparison of GraphOracle with Supervised SOTA Methods
B.1 Time Complexity Analysis
@@
GraphOracle’s overall time complexity consists of two parts: a one-time preprocessing cost and a per-query inference cost. Building the Relation–Dependency Graph (RDG) by scanning all triples once requires $\mathcal{O}(|F|),$ where $|F|$ is the total number of triples. For each query $(e_{q},r_{q},?)$, the $L_{R}$ layers of relation-level message passing on the RDG incur $\mathcal{O}(L_{R}\,|E_{R}|\,d),$ where $|E_{R}|$ is the number of edges in the RDG and $d$ is the hidden dimension; in practice $|E_{R}|\ll|F|$ and $L_{R}\leq 3$, so this cost is small. Subsequently, the $L_{E}$ layers of entity-level message passing propagate representations across an average branching factor $b$, costing $\mathcal{O}(L_{E}\,b\,d).$ With typical settings $L_{E}\leq 3$ and $b<30$, this yields sub-millisecond latency per query. Hence the end-to-end per-query complexity is $\mathcal{O}\bigl{(}L_{R}\,|E_{R}|\,d+L_{E}\,b\,d\bigr{)},$ while preprocessing remains $\mathcal{O}(|F|)$. Thanks to small constant depths and modest branching, GraphOracle achieves near-linear scalability and memory-efficient inference on large, heterogeneous knowledge graphs.
B.2 Relationship Dependency Graph (RDG) Construction Comparison
(a) KG1
(b) INGRAM
(c) ULTRA
(d) GraphOracle
(e) KG2
(f) INGRAM
(g) ULTRA
(h) GraphOracle
Dataset # Relation INGRAM ULTRA GraphOracle
NL-25 146 1610 2300 797
NL-50 150 1748 2526 861
NL-75 138 1626 2336 787
NL-100 99 892 1159 416
WK-25 67 598 947 256
WK-50 102 1130 2164 508
WK-75 77 732 1253 313
WK-100 103 1052 1695 460
FB-25 233 7172 10479 3501
FB-50 228 6294 9300 3135
FB-75 213 5042 7375 2524
FB-100 202 4058 5728 2017
WN_V1 9 48 40 37
WN_V2 10 76 76 55
WN_V3 11 94 85 68
WN_V4 9 70 61 54
FB_V1 180 1622 2416 712
FB_V2 200 2692 4050 1237
FB_V3 215 3398 5015 1640
FB_V4 219 4624 7036 2231
NL_V1 14 122 170 51
NL_V2 88 1574 2065 842
NL_V3 142 1942 2558 1017
NL_V4 76 1296 1657 744
(a) Knowledge Graph
(b) Relation Graph
Appendix C Statistics of Datasets
@@
Table 4: Statistics of the transductive KGs datasets. $Q_{\text{tra}}$, $Q_{\text{val}}$, $Q_{\text{tst}}$ are the query triplets used for reasoning.
@@
Dataset Reference # Entity # Relation $||\mathcal{E}|$$|Q_{\text{tra}}|$ $|Q_{\text{val}}|$ $|Q_{\text{tst}}|$ Supervised SOTA
WN18RR Dettmers et al. (2017) [8] 40.9k 11 65.1k 21.7k 3.0k 3.1k one-shot-subgraph [55]
FB15k237 Toutanova and Chen (2015) [36] 14.5k 237 204.1k 68.0k 17.5k 20.4k Adaprop [52]
NELL-995 Xiong et al.(2017) [44] 74.5k 200 112.2k 37.4k 543 2.8k Adaprop [52]
YAGO3-10 Suchanek et al.(2007) [32] 123.1k 37 809.2k 269.7k 5.0k 5.0k one-shot-subgraph [55]
Nell-V1 Teru et al. (2020) [34] 3.1k 14 5.5k 4.7k 0.4k 0.4k Adaprop [52]
Nell-V2 Teru et al. (2020) [34] 2.6k 88 10.1k 8.2k 90.9k 1.0k ULTRA [11]
Nell-V3 Teru et al. (2020) [34] 4.6k 142 20.1k 16.4k 1.9k 1.9k ULTRA [11]
Nell-V4 Teru et al. (2020) [34] 2.1k 76 9.3k 7.5k 0.9k 0.9k ULTRA [11]
WN-V1 Teru et al. (2020) [34] 2.7k 9 6.7k 5.4k 0.6k 0.6k Adaprop [52]
WN-V2 Teru et al. (2020) [34] 7.0k 10 20.0k 15.3k 1.8k 1.9k Adaprop [52]
WN-V3 Teru et al. (2020) [34] 12.1k 11 32.2k 25.9k 3.1k 3.2k Adaprop [52]
WN-V4 Teru et al. (2020) [34] 3.9k 9 9.8k 7.9k 0.9k 1.0k Adaprop [52]
FB-V1 Teru et al. (2020) [34] 1.6k 180 5.3k 4.2k 0.5k 0.5k ULTRA [11]
FB-V2 Teru et al. (2020) [34] 2.6k 200 12.1k 9.7k 1.2k 1.2k ULTRA [11]
FB-V3 Teru et al. (2020) [34] 3.7k 215 22.4k 18.0k 2.2k 2.2k ULTRA [11]
FB-V4 Teru et al. (2020) [34] 4.7k 219 33.9k 27.2k 3.4k 3.4k ULTRA [11]
Nell-25 Lee et al. (2023) [19] 5.2k 146 19.1k 17.6k 0.7k 0.7k ULTRA [11]
Nell-50 Lee et al. (2023) [19] 5.3k 150 19.3k 17.6k 0.9k 0.9k ULTRA [11]
Nell-75 Lee et al. (2023) [19] 3.3k 138 12.3k 11.1k 6.1k 6.1k ULTRA [11]
Nell-100 Lee et al. (2023) [19] 2.1k 99 9.4k 7.8k 0.8k 0.8k ULTRA [11]
WK-25 Lee et al. (2023) [19] 13.9k 67 44.1k 41.9k 1.1k 1.1k ULTRA [11]
WK-50 Lee et al. (2023) [19] 16.3k 102 88.9k 82.5k 3.2k 3.2k ULTRA [11]
WK-75 Lee et al. (2023) [19] 8.1k 77 31.0k 28.7k 1.1k 1.1k ULTRA [11]
WK-100 Lee et al. (2023) [19] 15.9k 103 58.9k 49.9k 4.5k 4.5k ULTRA [11]
FB-25 Lee et al. (2023) [19] 8.7k 233 103.0k 91.6k 5.7k 5.7k ULTRA [11]
FB-50 Lee et al. (2023) [19] 8.6k 228 93.1k 85.4k 3.9k 3.9k ULTRA [11]
FB-75 Lee et al. (2023) [19] 6.9k 213 69.0k 62.8k 3.1k 3.1k ULTRA [11]
FB-100 Lee et al. (2023) [19] 6.5k 202 67.5k 62.8k 2.3k 2.3k ULTRA [11]
PrimeKG Chandak et al. (2023) [3] 85.0k 14 3911.9k 3129.8k 391.2k 391.2k Adaprop [52]
Amazon-book Wang et al. (2019) [41] 3404.2k 40 3404.2k 3210.3k 98.0k 95.9k KUCNet [22]
GeoKG GeoNames Team (2025) [15] 2054.2k 681 2784.5k 2673.1k 55.7k 55.7k one-shot-subgraph [55]
Appendix D Training details
D.1 Comprehensive Training Details
Sequential multi-dataset schedule.
@@ D.1 Comprehensive Training Details
Dataset-specific hyper-parameters.
@@
Learning-rate decay and early stopping.
At epoch $\epsilon$ we apply $\alpha_{k}^{(t+1)}=\gamma_{k}\!\cdot\!\alpha_{k}^{(t)}$; training terminates when validation MRR has not improved for $\tau=10$ epochs.
@@
where $\mathcal{T}$ preserves (i) relation-dependency graph weights and (ii) shared layer norms, while re-initializing dataset-specific embeddings.
@@
Appendix E Cross-domain Dataset Processing Details
@@
E.1 Processing Details of Biomedical Domain
E.1.1 Dataset Creation
@@
E.1.2 External Information-Enrichment Dataset Creation
@@
E.2 Processing Detail of Recommendation Domain
@@ 35 PersistEntityMapping$(\mathcal{M}_{E})$;
Phase I (Relation Ontology Extension) introduces a formal extension of the relational schema $\mathcal{R}$ with a domain-specific "purchase" relation. Let $\bm{\Omega}=\{(r_{1},id_{1}),(r_{2},id_{2}),...,(r_{n},id_{n})\}$ represent the initial relation set. The algorithm derives $r_{max}=\max_{r\in\bm{\Omega}}(\text{id}(r))$ and establishes $\text{id}_{purchase}=r_{max}+1$, thus creating an extended relation ontology $\mathcal{R}^{\prime}=\bm{\Omega}\cup\{(\text{"purchase"},\text{id}_{purchase})\}$. This expansion facilitates the semantic representation of user-item interactions within the KG structure.
@@
E.3 Processing Detail of Geographic Datasets (GeoKG)
@@ 29 Build index maps $\pi_{\mathcal{E}},\pi_{\mathcal{R}}$ by ascending order of IDs;
E.3.1 Fully-Inductive Geographic KG Dataset Construction.
@@
Appendix F Complete Experimental Results
In this section, we report the comprehensive experimental results of our study. Table 8 presents the performance of GraphOracle on transductive benchmarks, while Table 9 and Table 10 summarize the results on entity-inductive and fully-inductive settings, respectively. Cross-domain evaluations are provided in Table 11, Table 12, and Table 13. Furthermore, Table 1 illustrates the enhanced performance of GraphOracle+ when external information is incorporated. Across all datasets and evaluation scenarios, GraphOracle consistently outperforms existing baselines by a notable margin, highlighting the robustness and effectiveness of our proposed approach.
@@ Appendix F Complete Experimental Results
Non-GNN ConvE 0.427 0.392 0.498 0.325 0.237 0.501 0.511 0.446 0.619 0.520 0.450 0.660
QuatE 0.480 0.440 0.551 0.350 0.256 0.538 0.533 0.466 0.643 0.379 0.301 0.534
RotatE 0.477 0.428 0.571 0.337 0.241 0.533 0.508 0.448 0.608 0.495 0.402 0.670
MINERVA 0.448 0.413 0.513 0.293 0.217 0.456 0.513 0.413 0.637
DRUM 0.486 0.425 0.586 0.343 0.255 0.516 0.532 0.460 0.662 0.531 0.453 0.676
AnyBURL 0.471 0.441 0.552 0.301 0.209 0.473 0.398 0.276 0.454 0.542 0.477 0.673
RNNLogic 0.483 0.446 0.558 0.344 0.252 0.530 0.416 0.363 0.478 0.554 0.509 0.622
RLogic 0.477 0.443 0.537 0.310 0.203 0.501 0.416 0.252 0.504 0.360 0.252 0.504
DuASE 0.489 0.448 0.569 0.329 0.235 0.519 0.423 0.372 0.592 0.473 0.387 0.628
GraphRulRL 0.483 0.446 0.541 0.385 0.314 0.575 0.425 0.278 0.527 0.432 0.354 0.517
GNNs CompGCN 0.479 0.443 0.546 0.355 0.264 0.535 0.463 0.383 0.596 0.421 0.392 0.577
NBFNet 0.551 0.497 0.666 0.415 0.321 0.599 0.525 0.451 0.639 0.550 0.479 0.686
RED-GNN 0.533 0.485 0.624 0.374 0.283 0.558 0.543 0.476 0.651 0.559 0.483 0.689
A*Net 0.549 0.495 0.659 0.411 0.321 0.586 0.549 0.486 0.652 0.563 0.498 0.686
AdaProp 0.562 0.499 0.671 0.417 0.331 0.585 0.554 0.493 0.655 0.573 0.510 0.685
ULTRA 0.480 0.479 0.614 0.368 0.339 0.564 0.509 0.462 0.660 0.557 0.531 0.710
one-shot-subgraph 0.567 0.514 0.666 0.304 0.223 0.454 0.547 0.485 0.651 0.606 0.540 0.721
GraphOracle 0.675 0.617 0.762 0.471 0.396 0.664 0.621 0.563 0.751 0.683 0.644 0.798
@@
GraphOracle 92.3 92.6 69.5 81.6 76.7 79.2 75.7 78.5 91.2 86.9 86.2 82.4
Table 10: Comparison of GraphOracle with other reasoning methods in relation&entity inductive settting. Best performance is indicated by the bold face numbers, and the underline means the second best. “H@1” and “H@10” are short for Hit@1 and Hit@10 (in percentage), respectively. “–” means unavailable results.
@@
GraphOracle 0.369 0.176 0.693 0.469 0.345 0.698 0.189 0.101 0.341 0.582 0.460 0.780
@@
GraphOracle 0.525 0.361 0.799 0.538 0.356 0.872 0.585 0.434 0.913 0.562 0.370 0.930

### v2

@@ 6 Conclusion
In this work, we introduced GraphOracle, a relation-centric foundation model for unifying reasoning across heterogeneous KGs. By converting KGs into RDG, our approach explicitly encodes compositional patterns among relations, yielding domain-invariant embeddings. Experiments on 60 diverse benchmarks showed consistent state-of-the-art performance, improving mean reciprocal rank by up to 16.8% over baselines with minimal adaptation. We also demonstrated that GraphOracle’s performance can be enhanced by integrating external information through GraphOracle+, which leverages foundation model embeddings for improved initialization. Ablation studies confirmed the essential contributions of the relation-dependency graph and multi-head attention components. These findings establish relation-dependency pre-training as a scalable approach toward universal KG reasoning, opening new avenues for cross-domain applications.
A Reproducibility and Code
To ensure reproducibility, we provide the complete source code of the GraphOracle framework in the supplementary materials.
B Limitations and Future Work
@@
C Time Complexity Analysis
@@
GraphOracle’s overall time complexity consists of two parts: a one‑time preprocessing cost and a per‑query inference cost. Building the Relation–Dependency Graph (RDG) by scanning all triples once requires $\mathcal{O}(|F|),$ where $|F|$ is the total number of triples. For each query $(e_{q},r_{q},?)$, the $L_{R}$ layers of relation‑level message passing on the RDG incur $\mathcal{O}(L_{R}\,|E_{R}|\,d),$ where $|E_{R}|$ is the number of edges in the RDG and $d$ is the hidden dimension; in practice $|E_{R}|\ll|F|$ and $L_{R}\leq 3$, so this cost is small. Subsequently, the $L_{E}$ layers of entity‑level message passing propagate representations across an average branching factor $b$, costing $\mathcal{O}(L_{E}\,b\,d).$ With typical settings $L_{E}\leq 3$ and $b<30$, this yields sub‑millisecond latency per query. Hence the end‑to‑end per‑query complexity is $\mathcal{O}\bigl(L_{R}\,|E_{R}|\,d+L_{E}\,b\,d\bigr),$ while preprocessing remains $\mathcal{O}(|F|)$. Thanks to small constant depths and modest branching, GraphOracle achieves near‑linear scalability and memory‑efficient inference on large, heterogeneous knowledge graphs.
D Statistics of Datasets
We choose the same datasets as ULTRA (galkin2024ultra), TRIX (zhang2025trix) and KG-ICL (cui2024promptkg). Details of the used Knowledge Graph datasets are given in Table 4 and Table 5. Furthermore, we first create three cross-domain datasets for cross-domain Knowledge Graph Reasoning, whose details can be found in Appendix F.
@@ D Statistics of Datasets
Table 4: Statistics of the KG datasets. $Q_{\text{tra}}$, $Q_{\text{val}}$, $Q_{\text{tst}}$ are the query triplets used for reasoning.
@@
Dataset Reference # Entity # Relation $|\mathcal{E}|$ $|Q_{\text{tra}}|$ $|Q_{\text{val}}|$ $|Q_{\text{tst}}|$ Supervised SOTA
WN18RR Dettmers et al. (2017) (dettmers2017convolutional) 40.9k 11 65.1k 21.7k 3.0k 3.1k one-shot-subgraph (one-shot-subgraph)
FB15k237 Toutanova and Chen (2015) (toutanova2015observed) 14.5k 237 204.1k 68.0k 17.5k 20.4k Adaprop (Adaprop)
NELL-995 Xiong et al.(2017) (xiong2017deeppath) 74.5k 200 112.2k 37.4k 543 2.8k Adaprop (Adaprop)
YAGO3-10 Suchanek et al.(2007) (suchanek2007yago) 123.1k 37 809.2k 269.7k 5.0k 5.0k one-shot-subgraph (one-shot-subgraph)
Nell-V1 Teru et al. (2020) (TER2020) 3.1k 14 5.5k 4.7k 0.4k 0.4k KG-ICL (cui2024promptkg)
Nell-V2 Teru et al. (2020) (TER2020) 2.6k 88 10.1k 8.2k 90.9k 1.0k KG-ICL (cui2024promptkg)
Nell-V3 Teru et al. (2020) (TER2020) 4.6k 142 20.1k 16.4k 1.9k 1.9k KG-ICL (cui2024promptkg)
Nell-V4 Teru et al. (2020) (TER2020) 2.1k 76 9.3k 7.5k 0.9k 0.9k KG-ICL (cui2024promptkg)
WN-V1 Teru et al. (2020) (TER2020) 2.7k 9 6.7k 5.4k 0.6k 0.6k KG-ICL (cui2024promptkg)
WN-V2 Teru et al. (2020) (TER2020) 7.0k 10 20.0k 15.3k 1.8k 1.9k KG-ICL (cui2024promptkg)
WN-V3 Teru et al. (2020) (TER2020) 12.1k 11 32.2k 25.9k 3.1k 3.2k KG-ICL (cui2024promptkg)
WN-V4 Teru et al. (2020) (TER2020) 3.9k 9 9.8k 7.9k 0.9k 1.0k KG-ICL (cui2024promptkg)
FB-V1 Teru et al. (2020) (TER2020) 1.6k 180 5.3k 4.2k 0.5k 0.5k KG-ICL (cui2024promptkg)
FB-V2 Teru et al. (2020) (TER2020) 2.6k 200 12.1k 9.7k 1.2k 1.2k KG-ICL (cui2024promptkg)
FB-V3 Teru et al. (2020) (TER2020) 3.7k 215 22.4k 18.0k 2.2k 2.2k KG-ICL (cui2024promptkg)
FB-V4 Teru et al. (2020) (TER2020) 4.7k 219 33.9k 27.2k 3.4k 3.4k KG-ICL (cui2024promptkg)
Nell-25 Lee et al. (2023) (lee2023ingram) 5.2k 146 19.1k 17.6k 0.7k 0.7k KG-ICL (cui2024promptkg)
Nell-50 Lee et al. (2023) (lee2023ingram) 5.3k 150 19.3k 17.6k 0.9k 0.9k KG-ICL (cui2024promptkg)
Nell-75 Lee et al. (2023) (lee2023ingram) 3.3k 138 12.3k 11.1k 6.1k 6.1k KG-ICL (cui2024promptkg)
Nell-100 Lee et al. (2023) (lee2023ingram) 2.1k 99 9.4k 7.8k 0.8k 0.8k KG-ICL (cui2024promptkg)
WK-25 Lee et al. (2023) (lee2023ingram) 13.9k 67 44.1k 41.9k 1.1k 1.1k KG-ICL (cui2024promptkg)
WK-50 Lee et al. (2023) (lee2023ingram) 16.3k 102 88.9k 82.5k 3.2k 3.2k KG-ICL (cui2024promptkg)
WK-75 Lee et al. (2023) (lee2023ingram) 8.1k 77 31.0k 28.7k 1.1k 1.1k KG-ICL (cui2024promptkg)
WK-100 Lee et al. (2023) (lee2023ingram) 15.9k 103 58.9k 49.9k 4.5k 4.5k KG-ICL (cui2024promptkg)
FB-25 Lee et al. (2023) (lee2023ingram) 8.7k 233 103.0k 91.6k 5.7k 5.7k KG-ICL (cui2024promptkg)
FB-50 Lee et al. (2023) (lee2023ingram) 8.6k 228 93.1k 85.4k 3.9k 3.9k KG-ICL (cui2024promptkg)
FB-75 Lee et al. (2023) (lee2023ingram) 6.9k 213 69.0k 62.8k 3.1k 3.1k KG-ICL (cui2024promptkg)
FB-100 Lee et al. (2023) (lee2023ingram) 6.5k 202 67.5k 62.8k 2.3k 2.3k KG-ICL (cui2024promptkg)
PrimeKG Chandak et al. (2023) (chandak2023building) 85.0k 14 3911.9k 3129.8k 391.2k 391.2k Adaprop (Adaprop)
Amazon-book Wang et al. (2019) (wang2019kgat) 3404.2k 40 3404.2k 3210.3k 98.0k 95.9k KUCNet (liu2024kucnet)
GeoKG GeoNames Team (2025) (geonames) 2054.2k 681 2784.5k 2673.1k 55.7k 55.7k one-shot-subgraph (one-shot-subgraph)
Table 5: Statistics of the KG datasets. $Q_{\text{tra}}$, $Q_{\text{val}}$, $Q_{\text{tst}}$ are the query triplets used for reasoning.
Dataset Reference # Entity # Relation $|Q_{\text{tra}}|$ $|Q_{\text{val}}|$ $|Q_{\text{tst}}|$ Supervised SOTA
CoDEx Small Dettmers et al. (2020) 2.0k 42 32.9k 1.8k 1.8k ULTRA (galkin2024ultra)
CoDEx Medium Dettmers et al. (2020) 17.1k 51 185.6k 10.3k 10.3k KG-ICL (cui2024promptkg)
CoDEx Large Dettmers et al. (2020) 78.0k 69 551.2k 30.6k 30.6k KG-ICL (cui2024promptkg)
WDsinger Lv et al. (2020) 10.3k 135 16.1k 2.2k 2.2k TRIX (zhang2025trix)
NELL23k Lv et al. (2020) 22.9k 200 25.4k 5.0k 5.0k KG-ICL (cui2024promptkg)
FB15k237_10 Lv et al. (2020) 11.5k 237 27.2k 15.6k 18.2k KG-ICL (cui2024promptkg)
FB15k237_20 Lv et al. (2020) 13.2k 237 54.4k 17.0k 20.0k KG-ICL (cui2024promptkg)
FB15k237_50 Lv et al. (2020) 14.1k 237 136.1k 17.4k 20.3k ULTRA (galkin2024ultra)
DBpedia100k Ding et al. (2018) 99.6k 470 597.6k 50.0k 50.0k TRIX (zhang2025trix)
AristoV4 Chen et al. (2021) 45.0k 1605 242.6k 20.0k 20.0k TRIX (zhang2025trix)
ConceptNet100k Malaviya et al. (2020) 78.3k 34 100.0k 1.2k 1.2k KG-ICL (cui2024promptkg)
Hetionet Himmelstein et al. (2017) 45.2k 24 2025.2k 112.5k 112.5k ULTRA (galkin2024ultra)
E Comprehensive Training Details
Sequential multi‑dataset schedule.
@@ E Comprehensive Training Details
Dataset‑specific hyper‑parameters.
@@
Learning‑rate decay and early stopping.
At epoch $\epsilon$ we apply $\alpha_{k}^{(t+1)}=\gamma_{k}\!\cdot\!\alpha_{k}^{(t)}$; training terminates when validation MRR has not improved for $10$ epochs.
@@
where $\mathcal{T}$ preserves (i) relation‑dependency graph weights and (ii) shared layer norms, while re‑initializing dataset‑specific embeddings.
@@
F Cross-domain Dataset Processing Details
@@
F.1 Processing Details of Biomedical Domain
[…]
Dataset Creation
@@ F.1 Processing Details of Biomedical Domain
External Information-Enrichment Dataset Creation
@@
F.2 Processing Detail of Recommendation Domain
@@ 31 $\Phi(\mathcal{E}_{\text{ord}}[j])\leftarrow j$;
Phase I (Relation Ontology Extension) introduces a formal extension of the relational schema $\mathcal{R}$ with a domain-specific ”purchase” relation. Let $\bm{\Omega}=\{(r_{1},id_{1}),(r_{2},id_{2}),...,(r_{n},id_{n})\}$ represent the initial relation set. The algorithm derives $r_{max}=\max_{r\in\bm{\Omega}}(\text{id}(r))$ and establishes $\text{id}_{purchase}=r_{max}+1$, thus creating an extended relation ontology $\mathcal{R}^{\prime}=\bm{\Omega}\cup\{(\text{"purchase"},\text{id}_{purchase})\}$. This expansion facilitates the semantic representation of user-item interactions within the KG structure.
@@
F.3 Processing Detail of Geographic Datasets (GeoKG)
MRR 0.013 0.032 0.019 0.021 0.075 0.084 0.093 0.425 0.486 0.473 0.493 0.473 0.528 0.539 0.639
H@1 0.028 0.064 0.039 0.042 0.088 0.097 0.104 0.403 0.443 0.428 0.450 0.428 0.486 0.493 0.552
H@10 0.046 0.088 0.048 0.053 0.103 0.125 0.176 0.534 0.566 0.564 0.573 0.586 0.627 0.654 0.793
@@ 29 Build index maps $\pi_{\mathcal{E}},\pi_{\mathcal{R}}$ by ascending order of IDs;
Fully-Inductive Geographic KG Dataset Construction.
@@
G Complete Experimental Results
In this section, we report the comprehensive experimental results of our study. Table 12 presents the performance of GraphOracle on transductive benchmarks, while Table 13, Table 14 and Table 15 summarize the results on entity-inductive and fully-inductive settings, respectively. Cross-domain evaluations are provided in Table 8, Table 10, and Table 11. Furthermore, Table 17 illustrates the enhanced performance of GraphOracle+ when external information is incorporated. Across all datasets and evaluation scenarios, GraphOracle consistently outperforms existing baselines by a notable margin, highlighting the robustness and effectiveness of our proposed approach.
@@ G Complete Experimental Results
Non-GNN ConvE 0.427 39.2 49.8 0.325 23.7 50.1 0.511 44.6 61.9 0.520 45.0 66.0
QuatE 0.480 44.0 55.1 0.350 25.6 53.8 0.533 46.6 64.3 0.379 30.1 53.4
RotatE 0.477 42.8 57.1 0.337 24.1 53.3 0.508 44.8 60.8 0.495 40.2 67.0
MINERVA 0.448 41.3 51.3 0.293 21.7 45.6 0.513 41.3 63.7
DRUM 0.486 42.5 58.6 0.343 25.5 51.6 0.532 46.0 66.2 0.531 45.3 67.6
AnyBURL 0.471 44.1 55.2 0.301 20.9 47.3 0.398 27.6 45.4 0.542 47.7 67.3
RNNLogic 0.483 44.6 55.8 0.344 25.2 53.0 0.416 36.3 47.8 0.554 50.9 62.2
RLogic 0.477 44.3 53.7 0.310 20.3 50.1 0.416 25.2 50.4 0.360 25.2 50.4
DuASE 0.489 44.8 56.9 0.329 23.5 51.9 0.423 37.2 59.2 0.473 38.7 62.8
GraphRulRL 0.483 44.6 54.1 0.385 31.4 57.5 0.425 27.8 52.7 0.432 35.4 51.7
GNNs CompGCN 0.479 44.3 54.6 0.355 26.4 53.5 0.463 38.3 59.6 0.421 39.2 57.7
NBFNet 0.551 49.7 66.6 0.415 32.1 59.9 0.525 45.1 63.9 0.550 47.9 68.6
RED-GNN 0.533 48.5 62.4 0.374 28.3 55.8 0.543 47.6 65.1 0.559 48.3 68.9
A*Net 0.549 49.5 65.9 0.411 32.1 58.6 0.549 48.6 65.2 0.563 49.8 68.6
AdaProp 0.562 49.9 67.1 0.417 33.1 58.5 0.554 49.3 65.5 0.573 51.0 68.5
ULTRA 0.480 47.9 61.4 0.368 33.9 56.4 0.509 46.2 66.0 0.557 53.1 71.0
one-shot-subgraph 0.567 51.4 66.6 0.304 22.3 45.4 0.547 48.5 65.1 0.606 54.0 72.1
TRIX 0.514 48.1 61.1 0.366 32.5 55.9 0.506 44.2 64.8 0.541 47.3 70.2
KG-ICL 0.536 49.6 63.7 0.376 32.7 53.8 0.534 46.7 67.2 0.545 47.4 68.8
GraphOracle 0.675 61.7 76.2 0.471 39.6 66.4 0.621 56.3 75.1 0.696 67.2 80.7
@@
TRIX 79.8 78.0 54.3 72.2 68.2 73.0 69.9 68.7 89.9 76.4 75.9 77.2
KG-ICL 82.7 78.7 62.6 74.9 70.0 76.8 70.4 70.6 99.5 83.5 79.9 80.2
GraphOracle 92.3 92.6 69.5 81.6 76.7 79.2 75.7 78.5 97.2 86.9 86.2 82.4
Table 14: Comparison of GraphOracle with other reasoning methods in fully-inductive setting. Best performance is indicated by the bold face numbers, and the underline means the second best. H@1” and H@10” are short for Hit@1 and Hit@10 (in percentage), respectively. –” means unavailable results.
@@
TRIX 0.188 0.093 0.290 0.368 0.254 0.513 0.166 0.078 0.313 0.300 0.354 0.401
KG-ICL 0.270 0.127 0.415 0.466 0.313 0.626 0.277 0.091 0.432 0.425 0.434 0.628
GraphOracle 0.417 0.192 0.711 0.469 0.345 0.698 0.362 0.101 0.498 0.582 0.460 0.780
Table 15: Comparison of GraphOracle with other reasoning methods in fully-inductive setting. Best performance is indicated by the bold face numbers, and the underline means the second best. H@1” and H@10” are short for Hit@1 and Hit@10 (in percentage), respectively. –” means unavailable results.
@@
TRIX 0.436 0.269 0.633 0.401 0.263 0.611 0.334 0.277 0.547 0.393 0.256 0.650
KG-ICL 0.499 0.307 0.719 0.458 0.274 0.664 0.384 0.291 0.598 0.434 0.279 0.694
GraphOracle 0.576 0.407 0.812 0.538 0.356 0.872 0.585 0.434 0.913 0.562 0.370 0.930
Table 16: Performance comparison among ULTRA, TRIX, KG-ICL, and GraphOracle across different datasets. Best results are in bold and second best are underlined.
Type Model ULTRA TRIX KG-ICL GraphOracle
MRR Hit@10 MRR Hit@10 MRR Hit@10 MRR Hit@10
Transductive CoDEx Small 0.490 0.686 0.484 0.676 0.479 0.662 0.512 0.697
CoDEx Medium 0.372 0.525 0.365 0.521 0.402 0.565 0.417 0.574
CoDEx Large 0.343 0.478 0.388 0.481 0.388 0.508 0.396 0.523
WDsinger 0.417 0.526 0.502 0.620 0.493 0.599 0.512 0.654
NELL23k 0.268 0.450 0.306 0.536 0.329 0.552 0.333 0.572
FB15k237_10 0.254 0.411 0.253 0.408 0.260 0.416 0.269 0.435
FB15k237_20 0.274 0.445 0.273 0.441 0.284 0.456 0.297 0.482
FB15k237_50 0.325 0.528 0.322 0.522 0.324 0.499 0.336 0.541
DBpedia100k 0.436 0.603 0.457 0.619 0.455 0.604 0.479 0.643
AristoV4 0.343 0.496 0.345 0.499 0.313 0.480 0.374 0.524
ConceptNet100k 0.310 0.529 0.340 0.564 0.371 0.584 0.386 0.602
Entity Inductive Hetionet 0.399 0.538 0.394 0.534 0.269 0.402 0.417 0.556
ILPC Small 0.303 0.453 0.310 0.455 0.316 0.473 0.339 0.497
ILPC Large 0.308 0.431 0.310 0.431 0.295 0.411 0.345 0.451
HM 1k 0.042 0.100 0.072 0.128 0.089 0.144 0.097 0.178
HM 3k 0.030 0.090 0.069 0.118 0.081 0.129 0.089 0.143
HM 5k 0.025 0.068 0.074 0.118 0.070 0.108 0.096 0.145
IndigoBM 0.432 0.639 0.436 0.645 0.440 0.641 0.483 0.697
Fully Inductive MT1 tax 0.330 0.459 0.397 0.508 0.411 0.521 0.491 0.568
MT1 health 0.380 0.467 0.376 0.457 0.387 0.479 0.405 0.501
MT2 org 0.104 0.170 0.098 0.162 0.100 0.171 0.132 0.193
MT2 sci 0.311 0.451 0.331 0.526 0.303 0.396 0.337 0.574
MT3 art 0.306 0.473 0.289 0.461 0.306 0.460 0.315 0.481
MT3 infra 0.657 0.807 0.672 0.810 0.676 0.808 0.697 0.829
MT4 sci 0.303 0.478 0.305 0.482 0.307 0.473 0.321 0.496
MT4 health 0.704 0.785 0.702 0.785 0.710 0.776 0.721 0.796
Metafam 0.997 1.000 0.997 1.000 1.000 1.000 1.000 1.000
FBNELL 0.481 0.661 0.478 0.655 0.516 0.699 0.523 0.732
NL-0 0.329 0.551 0.385 0.549 0.555 0.765 0.566 0.777
H Detail Design of GraphOracle+